OpenStudy (anonymous):

Can someone help me with. 1. Select two ordered pair points of your own. (Just pick any two points that seem relatively easy to work with.) Be sure to list these points for your instructor. Find the slope and show your work. 2. Take the points you have selected and write the equation of the line using the point-slope formula. Change the equation to slope-intercept form. Show the work. 3. Take the points you have selected and write the equation of the line using the slope-intercept formula. Show the work. 4. Compare and contrast the two methods of writing eq

OpenStudy (anonymous):

4. Compare and contrast the two methods of writing equations. List at least two similarities and two differences of using the two methods. Tell which process you like best and why.

OpenStudy (anonymous):

which two methods? personally i like solving out for slope intercept form, because you can easily find the interecpt, then draw the line depending on slope

OpenStudy (anonymous):

umm point-slope and slope-intercept form

OpenStudy (anonymous):

they arent really that different y2-y1= m(x2-x1) let (0,c) be the y intercept y2-0 = m(x2-b) y2=mx -mc let mc = b y2=mx2-b

OpenStudy (anonymous):

ok